The capital structure of Dog CC is current at 60% of equity (i.e., 40% of the debt). You find the debt issued by Dog CC has a pretax interest rate of 12%. You also find Dog CC issues its common stocks at $15.00 per share, with the current dividend of $1.00. You expect future dividends of Dogg ICC to grow by 4%.

what is DOG CC WACC if it pays taxes at 21%?
Answer #1

Hi

Here Pretax cost of debt kd = 12%

As per dividend discount model

Dividend growth rate g = 4%

Cost of equity ke = ?

Current Price of Share = Current Dividend*(1+g)/(ke-g)

15 = 1*(1+4%)/(ke-0.04)

15*ke - 0.6 = 1.04

ke = 1.64/15

ke = 10.93%

So WACC = D/V*(1-tax)*kd + E/V*ke

= 0.4*(1-0.21)*12 + 0.6*10.93

   = 3.79 +6.56

   =10.35%

Thanks